OpenStudy (bananas1234):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? 9, 40, 41 5, 10, 15 6, 9, 10 13, 35, 37 No idea on this one. :(

OpenStudy (butterflydreamer):

for this question, you want to think about pythagoras' theorem :) Pythagoras' theorem will only work on right angle triangles right? AND remember that the hypotenuse is always the longest side. So a^2 + b^2 = c^2 (where c is the hypotenuse) If you substitute in the values and this comes up as TRUE, then you know it's a right-angle triangle because pythagoras' theorem will apply ^_^ and ONLY works on right-angled triangles.

OpenStudy (bananas1234):

9^2 + 40^2 = 41^2? for the first one

OpenStudy (butterflydreamer):

yeeep that's exactly what you do :) So now just plug in "9^2 + 40^2" and find out what that equals to. Then find what 41^2 = If they're the SAME, then it's a right-angled triangle :D

OpenStudy (bananas1234):

ok, so 81 +1600 = 1681 1681 = 1681 This is true, so the answer would be A, the first one? @butterflydreamer

OpenStudy (butterflydreamer):

yaaaas :D!

OpenStudy (bananas1234):

Yay! thanks!!

OpenStudy (butterflydreamer):

no problemo!